40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 10 11:07:37 2011, rana, HowTo, SUS, Optical Lever Servo Tuning thoughts 
    Reply  Sun Mar 25 20:13:31 2012, rana, HowTo, SUS, Optical Lever Servo Tuning thoughts Untitled.png
Message ID: 6442     Entry time: Sun Mar 25 20:13:31 2012     In reply to: 5376
Author: rana 
Type: HowTo 
Category: SUS 
Subject: Optical Lever Servo Tuning thoughts 

To start the optical lever filter design, I looked into the noise on ITMY. It should be similar to the other arm cavity optics since they have the same whitening electronics.

The RED/BLUE are with loops open. The MAGENTA/CYAN with loops closed. Looks good; the bandwidth is a few Hz and there is not much peaking,

To figure out the contribution from the dark noise I misaligned the ITMY until the sum on the QPD went to zero. Then I took the spectra of the OL{1,2,3,4}_OUT signals (they all looked the same).

To normalize them properly I took OL4, multiplied it by 2 to account for the incoherent sum of 4 channels and then divided by the nominal SUM (which was 14685 counts). I've left the OL3 un-normalized to show the ratio.

From this plot it seems that the dark noise is not a problem at any frequency (no need to amplify for the new ADCs).

I'm going to use the open loop spectra to design the optimal feedback control. The file is saved as /users/rana/dtt/ITMY_OL-120325.xml

Attachment 1: Untitled.png  59 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-